give the quadratic equation whose root are 4 and 7

Answer by Horlaye(27)

Solution x = 4 or x = 7 x - 4 =0 or x - 7 =0 (x - 4)(x - 7) =0 x^2 - 7x - 4x + 28 =0 x^2 - 11x + 28 =0 hence the quadratic equation whose root are 4 and 7 = x^2 - 11x + 28 =0